# Gauge fields induced by curved spacetime

### v2

@@ Abstract
I found an extended duality (triality) between Dirac fermions in periodic spacetime metrics, nonrelativistic fermions in gauge fields (e.g., Harper-Hofstadter model), and in periodic scalar fields on a lattice (e.g., Aubry-André model). This indicates an unexpected equivalence between spacetime metrics, gauge fields, and scalar fields on the lattice, which is understood as different physical representations of the same mathematical object, the quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$. This quantum group is generated by the exponentiation of two canonical conjugate operators, namely a linear combination of position and momentum (periodic spacetime metrics), the two components of the gauge invariant momentum (gauge fields), position and momentum (periodic scalar fields). Hence, on a lattice, Dirac fermions in a periodic spacetime metric are equivalent to nonrelativistic fermions in a periodic scalar field after a proper canonical transformation. The three lattice Hamiltonians (periodic spacetime metric, Harper-Hofstadter, and Aubry-André) share the same properties, namely fractal phase diagrams, self-similarity, topological invariants, flat bands, and topological quantized current in the incommensurate regimes.
@@
In the light of the above, the result obtained here is surprising: On a discrete lattice[2], a periodic spacetime metric is equivalent to a gauge field and to a periodic scalar field. More precisely, I found that the Hamiltonian of a Dirac fermion in a curved spacetime[3, 4] described by a periodic spacetime metric is equivalent to the Hamiltonian of a nonrelativistic fermion in a gauge field (Harper-Hofstadter model[5, 6]) or in a periodic scalar field (Aubry-André model[7]) when regularized on the lattice. While the duality between nonrelativistic lattice fermions in gauge fields (Harper-Hofstadter) and periodic scalar fields (Aubry-André model) is well known, the connection with Dirac fermions in curved periodic spacetime on a lattice is unexpected: In essence, these three different models are all different physical representations of the same quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$, generated by the exponentiation of two canonical conjugate operators. These operators are the two components of the gauge invariant momentum (Harper-Hofstadter model), position and momentum (Aubry-André model), and a linear combination of position and momentum (periodic spacetime metrics). Hence, on a lattice, Dirac fermions in a periodic spacetime metric are obtained via a canonical transformation of nonrelativistic fermions in a periodic scalar field. This triality (i.e., nontrivial equivalence between three different models) mandates that all three models share the same properties, e.g., fractal phase diagrams, self-similarity, topological invariants, flat bands and topological quantized current in the incommensurate regimes, as I will shortly illustrate. Hamiltonians describing Dirac fermions in a curved spacetime regularized on a lattice are often considered in condensed matter[8, 9, 10, 11, 12, 13, 14]. These results may open new venues for studying analog gravity.
Nonrelativistic fermions on a 2D discrete lattice in a gauge field are described by the so-called Harper-Hofstadter Hamiltonian. To derive this Hamiltonian, one can start from charged fermions in a gauge field and then treat the discrete 2D lattice (described by a periodic potential) as a perturbation on the Landau levels (as in Ref. 15). Conversely, one can start with a so-called "tight-binding" model describing charged fermions in a periodic potential and treating the gauge field as a perturbation (as in Refs. 5, 6). These two approaches lead to the same result, namely
@@
where the two terms are the contributions of the gauge-invariant momentum $\hat{\boldsymbol{\pi}}=\hat{\mathbf{p}}-\hat{\mathbf{A}}$ regularized on the lattice, where $\hat{\mathbf{A}}=(-\omega_{y}\hat{y},\omega_{x}\hat{x})$ of a uniform field in Coulomb gauge. The two components of the gauge-invariant momentum are canonical conjugate operators $[\hat{\pi}_{x},\hat{\pi}_{y}]=\mathrm{i}\omega$ with $\omega=\omega_{x}+\omega_{y}$ equal to the flux of the field per unit cell.
@@
where the first term is the contribution of the scalar field $\propto\cos{\left(\omega{x}+\phi\right)}$ ), and the second term is the particle momentum regularized on the lattice.
@@
where $\hat{X},\hat{Y}$ are the canonical conjugate operators $[\hat{X},\hat{Y}]=\mathrm{i}\omega$. Thus, taking $\hat{X}=\hat{p}_{x}$ and $\hat{Y}=\hat{p}_{y}$ yields the Harper-Hofstadter model in Eq. 1; taking $\hat{X}=\omega\hat{x}+\phi$ and $\hat{Y}=\hat{p}$ yields the Aubry-André model in Eq. 2. The quantity $\omega=-\mathrm{i}[\hat{X},\hat{Y}]$ describes the flux of the gauge field per unit cell in the Harper-Hofstadter model and the wavenumber of the periodic scalar field in the Aubry-André model. The phase shift $\phi$ of the periodic scalar field in the Aubry-André model plays the same role as one of the components of the momentum in the Harper-Hofstadter model and, for this reason, can be regarded as a "synthetic" dimension[16]. In the quantum geometry language[17, 18, 19, 20, 21, 22, 23], the Hamiltonian Eq. 3 is $\mathcal{H}_{XY}=\hat{T}_{X}+\hat{T}_{X}^{\dagger}+\hat{T}_{Y}+\hat{T}_{Y}^{\dagger}$ where $\hat{T}_{X}=\mathrm{e}^{\mathrm{i}\hat{X}}$ and $\hat{T}_{Y}=\mathrm{e}^{\mathrm{i}\hat{Y}}$ are the so-called magnetic translation operators, satisfying the algebraic equations $\hat{T}_{X}\hat{T}_{Y}=\mathrm{e}^{\mathrm{i}\omega}\hat{T}_{Y}\hat{T}_{X}$, which are a quantum "deformation" of the usual algebra of spatial translations, which is commutative $\hat{T}_{X}\hat{T}_{Y}=\hat{T}_{Y}\hat{T}_{X}$. These operators $\hat{T}_{X}$ and $\hat{T}_{Y}$ coincide with two of the generators of the quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ obtained as a $q$-deformation of the enveloping algebra $U(\mathfrak{sl}_{2})$ of the Lie algebra $\mathfrak{sl}_{2}$, with $q^{2}=\mathrm{e}^{\mathrm{i}\omega}$.
However, the story does not necessarily end here: In principle, one can consider any set of conjugate variables $\hat{X},\hat{Y}$ in Eq. 3. An interesting choice is $\hat{X}=\frac{1}{2}(\omega\hat{x}+\phi)-\hat{p}$ and $\hat{Y}=\frac{1}{2}(\omega\hat{x}+\phi)+\hat{p}$ giving again $[\hat{X},\hat{Y}]=\mathrm{i}\omega$ and
@@
where the subscript stands for curved spacetime, as clarified hereafter. Regularizing on the lattice
@@
$\mathcal{H}=\sum\limits_{n}t_{n}\left(\psi_{n+1}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n}+\psi_{n}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n+1}\right).$ (11)
Now, consider the metric
@@
for $n\geq 1$ and with $C$ an arbitrary number defining the boundary value $\alpha_{1}=C$. In this metric $2t_{n}=\sqrt{{\alpha_{n}\alpha_{n+1}}}=|\cos{\left(\frac{1}{2}(\omega n+\phi)\right)}|$, and therefore the Hamiltonian $\mathcal{H}$ in Eq. 11 becomes equivalent to the Hamiltonian $\mathcal{H}_{\text{CS}}$ in Eq. 4 describing massless Dirac fermion in a periodic spacetime metric on a 1D lattice.
When the wavelength is commensurate with the lattice, i.e., when $\omega=2\pi p/q$ with $p,q$ coprimes, the hopping amplitudes are periodic. Conversely, when the wavelength is incommensurate, the hopping amplitudes are quasiperiodic. Note that the metric on the lattice $\alpha_{n}$ diverges for values of the phase $\omega m+\phi=\pm\pi$ on lattice sites where $m+n$ is even. The values of the phase where the metric is singular form a set $\{\phi=\pm\pi-\omega m\mod 2\pi,m\in\mathbb{Z}\}\subset[0,2\pi]$. In the commensurate case, the cardinality of the set is $q$ or $2q$ depending on whether $q$ is even or odd, i.e., there is a finite number of values ($q$ or $2q$) of the phase $\phi$ where the metric is singular. In the incommensurate case, the cardinality of the set is $\aleph_{0}$, i.e., there is an infinite (and countable) number of values of the phase where the metric is singular. Note that, even in cases where the metric is singular, the hopping amplitudes $t_{n}\propto\sqrt{\alpha_{n}\alpha_{n+1}}$ remain finite and well-defined. Note also that $\alpha_{n}$ is not, in general, periodic, in contrast with the hopping amplitudes $t_{n}$ that are always periodic in the commensurate regime or quasiperiodic in the incommensurate regime. In particular, one can show that $\alpha_{n}$ is periodic on the lattice with period $2q$ in the commensurate regime only for $q$ odd [^1] See Supplemental Material. However, for the sake of simplicity, I refer to the metric as being periodic or quasiperiodic as long as the hopping amplitudes on the lattice are periodic or quasiperiodic.
@@
Figure 2: Energy spectra of the Hamiltonian $\mathcal{H}_{\text{CS}}$ of a massless Dirac fermion in a periodic spacetime metric on a lattice of $N=210$ sites. (a) Energy spectra as a function of $p/q$ with periodic boundary conditions and $\phi=0$. (b) Energy spectra for $p/q=1/3$ as a function of the phase $\phi$ with open boundary conditions. (c) The metric $\alpha_{n}$ for $p/q=1/3$ and $\phi=0$ compared with $\cos{\left(\frac{1}{2}(\omega n+\phi)\right)}$ ) (dashed) as a function of the lattice site. (d) Wavefunctions of the two edge modes in the first band gap (from below) for $p/q=1/3$ and $\phi=0$. (e) Energy spectra for $\omega\approx 2\pi(\Phi-1)$ (with $\Phi$ the golden ratio), as a function of the phase $\phi$ with periodic boundary conditions on a lattice of $N=233$ sites. (f) The metric $\alpha_{n}$ for $\omega\approx 2\pi(\Phi-1)$ and $\phi=0$ as a function of the lattice site. (g) Wavefunctions of two bulk modes for $\omega\approx 2\pi(\Phi-1)$ and $\phi=0$.
@@
If the wavelength is commensurate to the lattice, the spectrum of the Hamiltonian $\mathcal{H}_{\text{CS}}$ in Eq. 4 has exactly $q$ energy bands spanning the phase $\phi$ and momentum $k$. Each of the open gaps $j$ is labeled by a nonzero topological invariant $c\in\mathbb{Z}$ given by the Chern number[27, 15, 28] $|c|<q/2$ satisfying the diophantine equation[29, 30, 31] $pc=j\mod q$. These nontrivial gaps exhibit $|c|$ topologically protected edge modes localized at the boundaries when considering a finite lattice of $N=mq$ sites with open boundary conditions, as it follows from the bulk-boundary correspondence[32].
For incommensurate wavelengths $\omega$, i.e., when $\omega/2\pi$ is an irrational number $\in\mathbb{R}-\mathbb{Q}$, the energy spectrum is a Cantor set[33], and the dispersion of the Bloch bands becomes flat[34, 35, 36]. Hence, the periodic spacetime metric induces a fractal phase diagram spanned by the wavelength $\omega$ with infinitely many nontrivial gaps, as in the Hofstadter butterfly[6].
This phase diagram shows a remarkable self-similarity property[21, 22]: It is invariant under the action of the duality transformation $p/q\to q/p,E\to\widetilde{E}$, with $\widetilde{E}$ given by some unknown function and with $p/q$ defined modulo 1 (or equivalently $\omega=2\pi p/q$ defined modulo $2\pi$). This property has been related to the notion of modular double[37] and, by extension, to the Langlands duality of quantum groups[22]. In the Harper-Hofstadter model, the two branches of the duality transformation $p/q\to q/p$ correspond to two opposite limits, obtained by considering the periodic potential as a perturbation on the Landau levels, or by considering the gauge field as a perturbation on the tight-binding model describing particles trapped in a periodic potential, as already noted in Ref. 15. In the Aubry-André model and in the curved spacetime model in Eq. 4, this duality transformation corresponds to switching the wavelength $\omega\to(2\pi)^{2}/\omega$ of the periodic scalar field and of the periodic spacetime metric, respectively. The spectrum is given by the roots of the characteristic polynomial expressed by the Chambers relation[38]
@@
where with abuse of notation, $\mathcal{H}_{\text{CS}}$, $\mathcal{H}_{\text{AA}}$ denote here the matrices expressing the corresponding Hamiltonians in the basis of plane waves $\mathrm{e}^{\mathrm{i}kn}$ on the lattice, with the function $f_{p/q}(E)$ given by a polynomial in $E$ (see Ref. 23). The spectrum is given by the roots of the characteristic polynomial $\mathcal{P}(E)=\det(\mathcal{H}_{\text{CS}}-E)=\det(\mathcal{H}_{\text{AA}}-E)$, where with abuse of notation, $\mathcal{H}_{\text{CS}}$, $\mathcal{H}_{\text{AA}}$ denote here the matrices expressing the Hamiltonians in the basis of plane waves $\mathrm{e}^{\mathrm{i}kn}$. This polynomial is given by the Chambers relation[38] $\mathcal{P}(E)=f_{p/q}(E)-2(-1)^{q}\left(\cos(qk)+\cos(q\phi)\right)$, with the function $f_{p/q}(E)$ given by a polynomial in $E$ (see Ref. 23). An analogous expression for $\det(\mathcal{H}_{\text{HH}}-E)$ is obtained by taking $k\to k_{x}$ and $\phi\to k_{y}$. Remarkably $f_{p/q}(E)=f_{q/p}(\widetilde{E})$ under the duality transformation $p/q\to q/p$[21, 22].

### v3

@@ Abstract
I found an extended duality (triality) between Dirac fermions in periodic spacetime metrics, nonrelativistic fermions in gauge fields (e.g., Harper-Hofstadter model), and in periodic scalar fields on a lattice (e.g., Aubry-André model). This indicates an unexpected equivalence between spacetime metrics, gauge fields, and scalar fields on the lattice, which is understood as different physical representations of the same mathematical object, the quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$. This quantum group is generated by the exponentiation of two canonical conjugate operators, namely a linear combination of position and momentum (periodic spacetime metrics), the two components of the gauge invariant momentum (gauge fields), position and momentum (periodic scalar fields). Hence, on a lattice, Dirac fermions in a periodic spacetime metric are equivalent to nonrelativistic fermions in a periodic scalar field after a proper canonical transformation. The three lattice Hamiltonians (periodic spacetime metric, Harper-Hofstadter, and Aubry-André) share the same properties, namely fractal phase diagrams, self-similarity, topological invariants, flat bands, and topologically quantized current in the incommensurate regimes. This work unveils an unexpected link between gravity and gauge fields, opens new venues for studying analog gravity, e.g., the Unruh effect and universe expansions/contractions, and hints at novel pathways to quantized gravity theories.
I Introduction
@@
In the light of the above, the result obtained here is surprising: On a discrete lattice [2], a periodic spacetime metric is equivalent to a gauge field and to a periodic scalar field. More precisely, I found that the Hamiltonian of a Dirac fermion in a curved spacetime [3, 4] described by a periodic spacetime metric is equivalent to the Hamiltonian of a nonrelativistic fermion in a gauge field (Harper-Hofstadter model [5, 6]) or in a periodic scalar field (Aubry-André model [7]) when regularized on the lattice. While the duality between nonrelativistic lattice fermions in gauge fields (Harper-Hofstadter) and periodic scalar fields (Aubry-André model) is well known [8, 9], the connection with Dirac fermions in curved periodic spacetime on a lattice is unexpected: In essence, these three different models are all different physical representations of the same quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$, generated by the exponentiation of two canonical conjugate operators. These operators are the two components of the gauge invariant momentum (Harper-Hofstadter model), position and momentum (Aubry-André model), and a linear combination of position and momentum (periodic spacetime metrics). Hence, on a lattice, Dirac fermions in a periodic spacetime metric are obtained via a canonical transformation of nonrelativistic fermions in a periodic scalar field. This triality (i.e., nontrivial equivalence between three different models) mandates that all three models share the same properties, e.g., fractal phase diagrams, self-similarity, topological invariants, flat bands and topological quantized current in the incommensurate regimes, as I will shortly illustrate. Hamiltonians describing Dirac fermions in a curved spacetime regularized on a lattice are often considered in condensed matter [10, 11, 12, 13, 14, 15, 16]. These results may open new venues for studying analog gravity, in particular the Unruh effect and universe expansions/contrations, and hint at novel ways to look at quantized gravity theories.
II Gauge fields and periodic scalar fields on a lattice
Nonrelativistic fermions on a 2D discrete lattice in a gauge field are described by the so-called Harper-Hofstadter Hamiltonian. To derive this Hamiltonian, one can start from charged fermions in a gauge field and then treat the discrete 2D lattice (described by a periodic potential) as a perturbation on the Landau levels (as in Ref. [17]). Conversely, one can start with a so-called "tight-binding" model describing charged fermions in a periodic potential and treating the gauge field as a perturbation (as in Refs. [5, 6]). These two approaches lead to the same result, namely
@@ II Gauge fields and periodic scalar fields on a lattice
where the two terms are the contributions of the gauge-invariant momentum $\hat{\boldsymbol{\pi}}=\hat{\mathbf{p}}-\hat{\mathbf{A}}$ regularized on the lattice, where $\hat{\mathbf{A}}=(-\omega_{y}\hat{y},\omega_{x}\hat{x})$ of a uniform field in Coulomb gauge. The two components of the gauge-invariant momentum are canonical conjugate operators $[\hat{\pi}_{x},\hat{\pi}_{y}]=\mathrm{i}\omega$ with $\omega=\omega_{x}+\omega_{y}$ equal to the flux of the field per unit cell. In the continuum limit at small momenta, Eq. 1 yields the Hamiltonian of charged particles in a uniform gauge field $\mathcal{H}_{\text{HH}}\sim(p_{x}+\omega_{y}y)^{2}+(p_{y}-\omega_{x}x)^{2}$ (up to a constant value), which in the Landau gauge is equivalent to the Hamiltonian of the quantum harmonic oscillator $\mathcal{H}_{\text{HH}}\sim p_{x}^{2}+\omega_{x}^{2}(x-k_{y}/\omega_{x})^{2}$ where the minimum of the potential depends on the eigenvalue of the momentum $k_{y}$. This is the celebrated Hamiltonian of the Landau levels which describes the physics of the integer quantum Hall effect of 2D electron systems, that exhibit topologically protected edge modes and quantized Hall conductance proportional to the topological Chern number (also known as the TKKN invariant introduced by Thouless, Kohomoto, Nightingale, and den Nijs [17]).
@@
where the first term is the contribution of the scalar field $\propto\cos{\left(\omega{x}+\phi\right)}$ ), and the second term is the particle momentum regularized on the lattice. In the continuum limit at small momenta and small frequencies $\omega$, Eq. 2 yields $\mathcal{H}_{\text{AA}}\sim p^{2}+\omega^{2}(x+\phi/\omega)^{2}$ (up to a constant value), which is again the quantum harmonic oscillator, now with the minimum of the potential depending on the phase shift $\phi$. The Hamiltonian in Eq. 2 can be generalized to $2J\cos{\left(\omega\hat{x}+\phi\right)}+2K\cos{\hat{p}}$ ). For almost every $\omega/2\pi,\phi/2\pi\in\mathbb{R}-\mathbb{Q}$ this Hamiltonian has exponentially localized eigenmodes in position space for $|J|>|K|$ and delocalized modes for $|J|<|K|$ with a phase transition, known as the Anderson localization transition, between the two regimes at $J=\pm K$ [18]. Hence, the Hamiltonian in Eq. 2 coincides with the Aubry-André Hamiltonian sitting right at the phase boundary between the localized and delocalized phases.
@@
where $\hat{X},\hat{Y}$ are the canonical conjugate operators $[\hat{X},\hat{Y}]=\mathrm{i}\omega$. Thus, taking $\hat{X}=\hat{p}_{x}$ and $\hat{Y}=\hat{p}_{y}$ yields the Harper-Hofstadter model in Eq. 1; taking $\hat{X}=\omega\hat{x}+\phi$ and $\hat{Y}=\hat{p}$ yields the Aubry-André model in Eq. 2. The quantity $\omega=-\mathrm{i}[\hat{X},\hat{Y}]$ describes the flux of the gauge field per unit cell in the Harper-Hofstadter model and the wavenumber of the periodic scalar field in the Aubry-André model. The phase shift $\phi$ of the periodic scalar field in the Aubry-André model plays the same role as one of the components of the momentum in the Harper-Hofstadter model and, for this reason, can be regarded as a "synthetic" dimension [19]. In the quantum geometry language [20, 21, 22, 23, 24, 25, 9], the Hamiltonian Eq. 3 is $\mathcal{H}_{XY}=\hat{T}_{X}+\hat{T}_{X}^{\dagger}+\hat{T}_{Y}+\hat{T}_{Y}^{\dagger}$ where $\hat{T}_{X}=e^{\mathrm{i}\hat{X}}$ and $\hat{T}_{Y}=e^{\mathrm{i}\hat{Y}}$ are the so-called magnetic translation operators, satisfying the algebraic equations $\hat{T}_{X}\hat{T}_{Y}=e^{\mathrm{i}\omega}\hat{T}_{Y}\hat{T}_{X}$, which are a quantum "deformation" of the usual algebra of spatial translations, which is commutative $\hat{T}_{X}\hat{T}_{Y}=\hat{T}_{Y}\hat{T}_{X}$. These operators $\hat{T}_{X}$ and $\hat{T}_{Y}$ coincide with two of the generators of the quantum group $\mathcal{U}_{q}(\mathfrak{sl}_{2})$ obtained as a $q$-deformation of the enveloping algebra $U(\mathfrak{sl}_{2})$ of the Lie algebra $\mathfrak{sl}_{2}$, with $q^{2}=e^{\mathrm{i}\omega}$.
However, the story does not necessarily end here: In principle, one can consider any set of conjugate variables $\hat{X},\hat{Y}$ in Eq. 3. An interesting choice is $\hat{X}=\frac{1}{2}(\omega\hat{x}+\phi)-\hat{p}$ and $\hat{Y}=\frac{1}{2}(\omega\hat{x}+\phi)+\hat{p}$ giving again $[\hat{X},\hat{Y}]=\mathrm{i}\omega$. Using the standard trigonometric sum-to-product identity, one gets
@@
where the subscript stands for curved spacetime, as clarified hereafter. In the continuum limit at small momenta and small frequencies, Eq. 4 yields $\mathcal{H}_{\text{CS}}\sim 2p^{2}+\frac{1}{2}\omega^{2}(x+\phi/\omega)^{2}$ (up to a constant value), which clearly describes a flat curvature geometry. Regularizing Eq. 4 on the lattice
@@ III Curved spacetime on a lattice
$\mathcal{H}=\sum\limits_{n}t_{n}\left(\psi_{n+1}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n}+\psi_{n}^{\dagger}\gamma_{0}\gamma^{1}\psi_{n+1}\right),$ (11)
which can be decomposed into two decoupled spin sectors in the Weyl representation $\gamma_{0}\gamma^{1}=-\sigma_{z}$. Now, consider the metric
@@
for $n\geq 1$ and with $C$ an arbitrary number defining the boundary value $\alpha_{1}=C$. In this metric $2t_{n}=\sqrt{{\alpha_{n}\alpha_{n+1}}}=|\cos{\left(\frac{1}{2}(\omega n+\phi)\right)}|$, and therefore each of the two decoupled spin sectors of the Hamiltonian $\mathcal{H}$ in Eq. 11 becomes equivalent to the Hamiltonian $\mathcal{H}_{\text{CS}}$ in Eq. 4 describing massless Dirac fermion in a periodic spacetime metric on a 1D lattice.
When the wavelength is commensurate with the lattice, i.e., when $\omega=2\pi p/q$ with $p,q$ coprimes, the hopping amplitudes are periodic. Conversely, when the wavelength is incommensurate, the hopping amplitudes are quasiperiodic. Note that the metric on the lattice $\alpha_{n}$ diverges for values of the phase $\omega m+\phi=\pm\pi$ on lattice sites where $m+n$ is even. The values of the phase where the metric diverges form a set $\{\phi=\pm\pi-\omega m\mod 2\pi,m\in\mathbb{Z}\}\subset[0,2\pi]$. In the commensurate case, the cardinality of the set is $q$ or $2q$ depending on whether $q$ is even or odd, i.e., there is a finite number of values ($q$ or $2q$) of the phase $\phi$ where the metric diverges. In the incommensurate case, the cardinality of the set is $\aleph_{0}$, i.e., there is an infinite (and countable) number of values of the phase where the metric diverges. Note that, even in cases where the metric diverges, the hopping amplitudes $t_{n}\propto\sqrt{\alpha_{n}\alpha_{n+1}}$ remain finite and well-defined, since in the limit $\phi\to\pm\pi-\omega m$, the quantity $\alpha_{n}\alpha_{n+1}$ converges on the whole lattice to a finite positive value or to zero, e.g., $t_{m}=0$. Note also that $\alpha_{n}$ is not, in general, periodic, in contrast with the hopping amplitudes $t_{n}$ that are always periodic in the commensurate regime or quasiperiodic in the incommensurate regime. In particular, one can show that $\alpha_{n}$ is periodic on the lattice with period $2q$ in the commensurate regime only for $q$ odd (see Appendix). However, for the sake of simplicity, I refer to the metric as being periodic or quasiperiodic as long as the hopping amplitudes on the lattice are periodic or quasiperiodic.
The lattice Hamiltonian $\mathcal{H}_{\text{CS}}$ describes fermions localized on a discrete set of points on a 1D line by the effect of, e.g., a scalar potential, and affected by the presence of a spacetime curvature in Eq. 7. Moreover, $\mathcal{H}_{\text{CS}}$ can be effectively reproduced in condensed matter by engineering a 1D lattice of fermions localized in ionic crystals or in optical lattices with variable hopping amplitudes.
@@
IV Topology, fractality, flat bands, and quantum geometry
@@
The topological properties of the models are encoded in a topological invariant [17, 27, 28], the first Chern number, which is an integer defined as $c=\frac{1}{2\pi}\sum_{n}^{\text{N}}\int_{T^{2}}\Omega_{n}(\boldsymbol{\kappa})\mathrm{d}^{2}\boldsymbol{\kappa}$, where $N$ is the number of occupied fermionic levels, $\Omega_{n}(\boldsymbol{\kappa})=\nabla_{\boldsymbol{\kappa}}\mathcal{A}_{n}(\boldsymbol{\kappa})$ the Berry curvature corresponding to the Berry connection $\mathcal{A}_{n}(\boldsymbol{\kappa})=\mathrm{i}\bra{u_{n}(\boldsymbol{\kappa})}\nabla_{\boldsymbol{\kappa}}\ket{u_{n}(\boldsymbol{\kappa})}$ of the eigenvectors |$\ket{u_{n}(\boldsymbol{\kappa})}$ (see, e.g., Ref. [29]). Here, the quantum numbers $\boldsymbol{\kappa}=(k_{x},k_{y})$ in the language of the Hamiltonian $\mathcal{H}_{\text{HH}}$, and $\boldsymbol{\kappa}=(\phi,k)$ in the language of the Hamiltonians $\mathcal{H}_{\text{AA}}$ and $\mathcal{H}_{\text{CS}}$, and the integral is over the whole space spanned by $\boldsymbol{\kappa}$, which is a torus $T^{2}$ (i.e., the Brillouin zone) due to the periodicity of the spectra with respect to $\kappa_{x}$ and $\kappa_{y}$. Due to the bulk-boundary correspondence [30, 31, 32], a number $c$ of topologically protected edge modes appear at the edges of a topologically nontrivial phase $c\neq 0$.
If the wavelength is commensurate to the lattice, the spectrum of the Hamiltonian $\mathcal{H}_{\text{CS}}$ in Eq. 4 has exactly $q$ energy levels spanning the phase $\phi$ and momentum $k$. Each of the open gaps $j$ is labeled by a nonzero topological invariant $c\in\mathbb{Z}$ given by the Chern number [27, 17] $|c|<q/2$ satisfying the diophantine equation [33, 34, 35] $pc=j\mod q$. These nontrivial gaps exhibit $|c|$ topologically protected edge modes localized at the boundaries when considering a finite lattice of $N=mq$ sites with open boundary conditions, as it follows from the bulk-boundary correspondence [30].
For incommensurate wavelengths $\omega$, i.e., when $\omega/2\pi$ is an irrational number $\in\mathbb{R}-\mathbb{Q}$, the energy spectrum is a Cantor set [36], and the dispersion of the Bloch bands becomes flat [37, 38, 39]. Hence, the periodic spacetime metric induces a fractal phase diagram spanned by the wavelength $\omega$ with infinitely many nontrivial gaps, as in the Hofstadter butterfly [6]. For almost every $\omega/2\pi,\phi/2\pi\in\mathbb{R}-\mathbb{Q}$, this regime coincides with the phase boundary (Anderson localization transition) between the localized and delocalized phase in the Aubry-André model in Eq. 2 [18].
This phase diagram shows a remarkable self-similarity property [24, 25]: It is invariant under the action of the duality transformation $p/q\to q/p,E\to\widetilde{E}$, with $\widetilde{E}$ given by some unknown function and with $p/q$ defined modulo 1 (or equivalently $\omega=2\pi p/q$ defined modulo $2\pi$). This property has been related to the notion of modular double [40] and, by extension, to the Langlands duality of quantum groups [25]. In the Harper-Hofstadter model, the two branches of the duality transformation $p/q\to q/p$ correspond to two opposite limits, obtained by considering the periodic potential as a perturbation on the Landau levels, or by considering the gauge field as a perturbation on the tight-binding model describing particles trapped in a periodic potential, as already noted in Ref. [17]. In the Aubry-André model and in the curved spacetime model in Eq. 4, this duality transformation corresponds to switching the wavelength $\omega\to(2\pi)^{2}/\omega$ of the periodic scalar field and of the periodic spacetime metric, respectively. The spectrum is given by the roots of the characteristic polynomial expressed by the Chambers relation [41]
@@
where with abuse of notation, $\mathcal{H}_{\text{CS}}$, $\mathcal{H}_{\text{AA}}$ denote here the matrices expressing the corresponding Hamiltonians in the basis of plane waves $e^{\mathrm{i}kn}$ on the lattice, with the function $f_{p/q}(E)$ given by a polynomial in $E$ (see Ref. [9]). This polynomial is given by the Chambers relation [41] $\mathcal{P}(E)=f_{p/q}(E)-2(-1)^{q}\left(\cos(qk)+\cos(q\phi)\right)$, with the function $f_{p/q}(E)$ given by a polynomial in $E$ (see Ref. [9]). An analogous expression for $\det(\mathcal{H}_{\text{HH}}-E)$ is obtained by taking $k\to k_{x}$ and $\phi\to k_{y}$. Remarkably $f_{p/q}(E)=f_{q/p}(\widetilde{E})$ under the duality transformation $p/q\to q/p$ [24, 25].